OpenStudy (anonymous):

Hello i am having trouble understanding what a logarithm is specifically, any help would be greatly appreciated

OpenStudy (anonymous):

4 and4 is 8, right?

OpenStudy (anonymous):

4 = 2^2 and 8 = 2^3, agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and 16 is 2^4

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

2^2 * 2^2 = 2^4, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What would 2^2 *2^2*2^2 be?

OpenStudy (anonymous):

2^6

OpenStudy (anonymous):

Which u got by adding the exponents, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's what a logarithm does, converts multiplication into addition.

OpenStudy (anonymous):

In this case u would say the base is 2.

OpenStudy (anonymous):

ok, that somewhat clarifys it, thank you

OpenStudy (anonymous):

Somewhat?

OpenStudy (anonymous):

its just that i dont understand WHAT a logrithim is, what you told me was how a logrithim is a base just as it is the exponents being added togeather

OpenStudy (anonymous):

log to base a of y MEANS y = a^x

OpenStudy (anonymous):

Is that clearer?

OpenStudy (anonymous):

yes, much

OpenStudy (anonymous):

x = log_a y is same as y = a^x

OpenStudy (anonymous):

4 = log_2 16 and 16 = 2^4

OpenStudy (anonymous):

Sure u got it?

OpenStudy (anonymous):

When the base is e, you say ln instead of log and call them "natural logarithms"

OpenStudy (anonymous):

ok! wow that really hlped thank you!

OpenStudy (anonymous):

if u were to ask the question. 2 to cubed equals what? u would have 2^3=y. y being what we are solving for. so 2^3=2*2*2=8 so the answer we are looking for is 8 from this we can get a general equation for values of 2 as we change the exponent 2^x=y so we change x to say 3 0 or 4 and we solve 2^4=16 so generally we are asking the question a number(like 2) to a power that we cahnge(or the x variable) equals what(the y variable). or a^x=y Now for logs we are asking a different question. suppose i said 8 is 2 to what power? then u would have 8=2^y and we put y here because it is what we are solving for of course we know this is 3 since we did it earlier but suppose i asked the question: now from this we can get a general equation to solve x=2^y as we change x we solve for the y values the even more general statement would b a number to what power equals the number we change or mathematically x=a^y now we take this equation and solve for y because we like it on 1 side by itself. but there is a problem here. how do we get the y by itself? so what we do is we create a log system which the opposite of the exponential system. thus all the rules of exponents and logs are reversed. so what we do is we take the log of each side. the base being a (or the constant) log(base a)(x)=log(base a)a^y log(base a)(x)=y now the important thing here is that log(base a)(x)=y and x=a^y is the same exact equation and asks the same exact question